# Astrometry meets Pulsar Timing Arrays: Synergies for Gravitational Wave Detection

### v1

@@ 2 Overlap functions for astrometry and pulsar timing arrays
In this section we obtain expressions for astrometric overlap reduction functions (ORF) and their cross-correlations with pulsar timing arrays. We include the possible presence of a dipolar SGWB anisotropy, which can be instrumental for distinguishing cosmological from astrophysical backgrounds. We investigate the crucial role of the position of the monitored objects for determining the sensitivity of the system to the SGWB properties. Our covariant, analytical formulas are convenient and simple to use for Fisher forecast analysis.
@@ 2.1 Astrometry overlap functions
where $f$ is the GW frequency, and ${\bf p}$ the unit vector indicating its direction. ${\bf e}_{ij}^{\lambda}({\bf p})$ are real polarization tensors in the basis $\lambda=(+,\times)$. We impose $h_{\lambda}(-f,{\bf p})\,=\,h^{*}_{\lambda}(f,{\bf p})$. A GW passing between a star and the Earth causes a distortion $\delta n^{i}$ on the star direction, ${\bf n}$, at time $t$, given by the formula (see e.g. [25])
$\delta n_{i}(t,{\bf n})\,=\,{\cal R}_{ikl}({\bf n},{\bf p})\,h_{kl}(t,0)$ (2.3)
${\cal R}_{ikl}({\bf n},{\bf p})\,=\,\frac{n_{k}}{2}\left[\frac{(n_{i}+p_{i})n_{l}}{1+{\bf n}\cdot{\bf p}}-\delta_{il}\right]$ (2.4)
@@
In order to characterize a GW signal, we focus on correlation functions of the distortions of star positions. The two-point function of the GW Fourier mode is expressed as
$\langle h_{\lambda_{1}}(f_{1},{\bf p}_{1})h^{*}_{\lambda_{2}}(f_{2},{\bf p}_{2})\rangle\,=\,\frac{3H_{0}^{2}}{4\pi^{2}}\,\frac{\Omega_{\rm GW}(f_{1},{\bf p}_{1})}{f_{1}^{3}}\,\delta(f_{1}-f_{2})\,\frac{\delta^{(2)}({\bf p}_{1}-{\bf p}_{2})}{4\pi}$ (2.5)
where $\Omega_{\rm GW}(f,{\bf p})$ is the quantity customarily used for characterizing the GW energy density per log frequency interval (see e.g. [94]). The spectral energy density parameter $\Omega_{\rm GW}$ is related to the GW intensity, $I$, through the relation
@@
[…]
@@
The tensors $H^{(0,1)}_{ij}$ are the analog of the PTA ORF relative to the monopolar and dipolar kinematic components of the SGWB. To compute these quantities, we introduce the projection tensor
@@
The integrals (2.9), (2.10), can be performed with a standard methods of contour integration used in the context of PTA physics, see e.g. [98]. For the case of the monopole, the ORF is associated with the integral in eq (2.9). We consider the combination
$y$ $=$ $\frac{1-{\bf n}\cdot{\bf q}}{2}\,=\,\frac{1-\cos{\zeta}}{2}\,.$ (2.12)
@@
a result equivalent to what was found in [25], although expressed slightly more compactly. As mentioned above, we obtained this expression by evaluating the integral (2.9) using contour integration. The expression (2.13) is the astrometry equivalent of the Hellings-Downs function, the PTA ORF for the SGWB monopole. In the case of astrometry, the value of (2.13) depends on the angular separation ${\bf n}\cdot{\bf q}$ between stars, as well as more specifically on the star directions $n_{i}$ and $q_{j}$.
@@
${\bf A}$ $=$ ${\bf n}\times{\bf q}\hskip 14.22636pt,\hskip 14.22636pt{\bf B}\,=\,{\bf n}\times{\bf A}\hskip 14.22636pt,\hskip 14.22636pt{\bf C}\,=\,-{\bf q}\times{\bf A}\,.$ (2.14)
@@
[…]
@@
Our analytical, covariant expressions (2.13) and (2.15) for the ORF demonstrate that the astrometry sensitivity to the SGWB depend not only on the angle among stars, $\zeta$, but also on their position in the sky, and on the velocity vector ${\bf v}$ among frames.
The compact expressions (2.13) and (2.15) are particularly suitable to visualise patterns of sensitivity on the celestial sphere, and to further explore geometrical features of physically relevant quantities in the context we are examining. We plot in Figure 1 the combinations
@@
Figure 1: The quantities ${\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]$ and ${\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]$, associated with response to a SGWB depending on stars positions. (See the main text for our notation.) Motivated by CMB, we choose the kinematic dipole direction ${\bf v}$ (red star) at $(l,b)=(264^{\circ},48^{\circ})$ in galactic coordinates. Each panel shows a different choice of ${\bf n}$, while the stars ${\bf q}$ take the position of each pixel of the map. Upper left panel: ${\bf n}$ pointing towards $(l,b)=(0,0)$. Upper right panel: ${\bf n}$ pointing towards $-{\bf v}$. Lower left panel: ${\bf n}$ at $(l,b)=(0,0)$. Lower right panel: ${\bf n}$ pointing towards the direction $(l,b)=(270.21^{\circ},-75.45^{\circ})$.
@@ 2.2 Cross-correlated overlap functions of astrometry and PTA
[…]
@@
$K^{(0)}_{i}({\bf n},{\bf x})$ $=$ $\int d^{2}\Omega_{{\bf p}}\,{\cal R}_{ikl}({\bf n},{\bf p})\,P_{klrs}\,\frac{x^{r}x^{s}}{1+{\bf x}\cdot{\bf p}}$ (2.21)
$K^{(1)}_{i}({\bf n},{\bf x},{\bf v})$ $=$ $\int d^{2}\Omega_{{\bf p}}\,({\bf p}\cdot{\bf v})\,\,{\cal R}_{ikl}({\bf n},{\bf p})\,P_{klrs}\,\frac{x^{r}x^{s}}{1+{\bf x}\cdot{\bf p}}$ (2.22)
@@
$K^{(0)}_{i}({\bf n},{\bf x})$ $=$ $\frac{16\pi}{3}\frac{(1-2y)\,n_{i}-x_{i}}{4y(1-y)}\,\left(2y-2y^{2}+3y^{2}\ln(y)\right)$ (2.24)
@@
We call $\zeta_{sp}$ the angle between star and pulsar directions, $\cos\zeta_{sp}={{\bf n}}\cdot{\bf{x}}$. We represent a simple example in the upper left panel of Figure 3 with $\bf{n}$ pointing towards $(l,b)=(0,0)$; the roots of the function $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$ are at $\zeta_{sp}=0,\pi$ and $\zeta_{sp}\approx 86.14^{\circ}$, while its local maxima are located at $\zeta_{sp}\approx 37.13^{\circ}$ and $\zeta_{sp}\approx 132.195^{\circ}$, and are represented in the left panel of Figure 4. The lower panel of Figure 3 shows two scenarios for the general expression for the function $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ (see also Appendix B), which depends on the angle between star and pulsar and the angle between the velocity and pulsars given by $vx={\bf v}\cdot{\bf{x}}$; on the left we choose $\bf{n}$ at $(l,b)=(0,0)$, and on the right, $\bf{n}$ is at position $(l,b)=(270.21^{\circ},-75.45^{\circ})$ (a direction chosen randomly). The sensitivity patterns are complex. However, in the specific case of the star direction $\bf{n}$ pointing towards the direction $-{\bf v}$ we obtain a simple formula, depending only on the angle between star and pulsar. This last case is shown on the right upper panel of Figure 3, where the bluest, less sensitive regions showing the positions of objects at relative angle $\zeta_{sp}=0,\pi$ and $39.82^{\circ},101.37^{\circ}$, corresponding to minimal sensitivity to dipole. The red region shows the positions of the pulsars where the function $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ reaches its maximum at $\zeta_{sp}\approx 142.29^{\circ}$. There are also local maxima at $\zeta_{sp}\approx 17.97^{\circ},71.88^{\circ}$, which are more apparent on the right panel of Figure 4. Notice that the quantities ${\bf K}_{0,1}$ corresponding to the synergetic ORF vanish in the concident limit of the star aligned with the pulsar direction. One may understand this as follows – the star deflection always lies in the plane perpendicular to the line of sight. On the contrary, the pulsar timing residual arises from the change in length along the line of sight. Thus for pulsar and star in same direction the two effects are perpendicular to each other.
@@
Figure 3: The quantities $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$, and $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ response to stars and pulsars positions. The dipole direction ${\bf v}$ (red star) is chosen in the direction $(l,b)=(264^{\circ},48^{\circ})$ in galactic coordinates. Each panel shows a different choice of ${\bf n}$, while the pulsars positions ${\bf{x}}$ scan over each pixel of the map. Upper left panel: ${\bf n}$ towards $(l,b)=(0,0)$. Upper right panel: ${\bf n}$ towards $-{\bf v}$. Lower left panel: ${\bf n}$ at $(l,b)=(0,0)$. Lower right panel: ${\bf n}$ towards $(l,b)=(270.21^{\circ},-75.45^{\circ})$.
@@
Figure 4: Overlap functions response in terms of the angle between the stars and pulsars. Left: Angular dependence for $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$. Right: The quantities $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ when choosing ${\bf n}=-{\bf v}$. .
3 Optimal estimators, and Fisher forecasts
@@
In section 3.1 we determine the general optimal estimators to detect the SGWB properties, and show how they depend on the star through the ORF determined in Section 2. Based on such general estimators, we build a Gaussian likelihood and derive concise expressions for the Fisher matrix, whose structure depend in a transparent way on the geometrical properties of the systems we consider. In appropriate limits and special cases, the Fisher matrix obeys interesting scaling relations which make our analysis particularly simple. In general though, numerical work is needed to evaluate at what extent the synergy between astrometry and PTA improve the sensitivity to SGWB properties. We carry out such analysis in sections 3.2, 3.2.3, making use of the weak signal approximation for astrometry to simplify the numerical calculations for a large number of stars.
3.1 Building the optimal estimator: the case of astrometry
We start by discussing the case of astrometry only. We expand the deflection correlation of eq (2.8) in a basis with coefficients $p_{n}$,
$\langle{\delta n_{a}^{i}\,\delta n_{b}^{j}}\rangle\equiv{\boldsymbol{C}}=\sum_{n=1,2}p_{n}H_{ab,n}^{ij}+N_{ab}^{ij}=\sum_{n=1,2}p_{n}{\boldsymbol{H}}_{n}+{\boldsymbol{N}}\,.$ (3.1)
The indexes $a,b$ identify the two stars. $i,j$ the three-dimensional vector components, and the index $n$ runs over $1,2$ indicating monopolar and dipolar contributions to the signal. We include the effects of noise controlled by the noise matrix $N_{ab}^{ij}$. The quantities $p_{n}$ and $H_{ab,n}$ are controlled by the properties of the SGWB, as well as the astrometry ORF. See equation (2.8) and the discussion that follows for the definitions. At this stage, we do not need to specify whether we work either in time or in frequency domain. Our aim is to determine the quantities $p_{0,1}$, and extract from this information the properties of the SGWB: its monopole amplitude $\Omega_{\rm GW}$ as well as the value of the parameter $\beta$ controlling kinematic anisotropies.
Before proceeding to build the optimal estimators, we notice that the matrix $C$ can not be directly be used as a covariance matrix, since it is singular. This stems from the fact that the three Cartesian deflection components are not independent since observations get projected on to the two-dimensional celestial sphere. Hence, the star deflections can be entirely described in terms of two angles $\delta\theta$ and $\delta\phi$ [28]. Let $P$ be the matrix that converts the deflections from 3D Cartesian to 2D polar coordinates, i.e.

### v2

@@ 2 Overlap functions for astrometry and pulsar timing arrays
The overlap reduction function (ORF) of a pair of detectors accounts for how the sensitivity to the SGWB changes due to the relative positions and orientations of the detectors. The cross-correlation of the signal across the two detectors can be written entirely in terms of the ORF and the underlying SGWB spectral density. Schematically, one can write
$\langle{s_{I}(f)s_{J}(f)}\rangle\propto\gamma_{IJ}(f)I(f)\,$, (2.1)
where $s_{I,J}$ denotes the signal at detector $I,J$, $\gamma_{IJ}$ the overlap reduction function and $I(f)$ the SGWB intensity. In this section we obtain expressions for the astrometric ORFs and their cross-correlations with pulsar timing arrays. We include the possible presence of a dipolar SGWB anisotropy, which can be instrumental for distinguishing cosmological from astrophysical backgrounds. We investigate the crucial role of the position of the monitored objects for determining the sensitivity of the system to the SGWB properties. Our covariant, analytical formulas are convenient and simple to use for Fisher forecast analysis.
@@ 2.1 Astrometry overlap functions
$h_{ij}(t,\vec{x})\,=\,\sum_{\lambda}\,\int_{-\infty}^{+\infty}df\,\int d^{2}{\bf p}\,{e}^{-2\pi if\,{\bf p}\cdot\vec{x}}\,e^{2\pi ift}\,{\bf e}_{ij}^{\lambda}({\bf p})\,h_{\lambda}(f,{\bf p})\,$, (2.3)
where $f$ is the GW frequency, and $\bf p$} the unit vector indicating its direction. ${\bf e}_{ij}^{\lambda}({\bf p})$ are real polarization tensors in the basis $\lambda=(+,\times)$. Notice that we take the flat space limit of the GW evolution, hence it satisfies a flat space d’Alembertian equation. We impose $h_{\lambda}(-f,{\bf p})\,=\,h^{*}_{\lambda}(f,{\bf p})$. A GW passing between a star and the Earth causes a distortion $\delta n^{i}$ on the star direction, $\bf n$}, at time $t$, given by the formula (see e.g. [25])
$\delta n_{i}(t,{\bf n})\,=\,{\cal R}_{ikl}({\bf n},{\bf p})\,h_{kl}(t,0)\,$, (2.4)
@@
In order to characterise the stochastic GW signal, we assume it to have a Gaussian distribution with zero mean.. Thus, the SGWB can be entirely characterised by its two-point function, expressed as
$\langle h_{\lambda_{1}}(f_{1},{\bf p}_{1})h^{*}_{\lambda_{2}}(f_{2},{\bf p}_{2})\rangle\,=\,\frac{3H_{0}^{2}}{4\pi^{2}}\,\frac{\Omega_{\rm GW}(f_{1},{\bf p}_{1})}{f_{1}^{3}}\,\delta_{\lambda_{1}\lambda_{2}}\,\delta(f_{1}-f_{2})\,\frac{\delta^{(2)}({\bf p}_{1}-{\bf p}_{2})}{4\pi}\,$, (2.6)
where the angle brackets denote an ensemble average, over all possible realisations of the SGWB and $\Omega_{\rm GW}(f,{\bf p})$ is the quantity customarily used for characterizing the GW energy density per log frequency interval (see e.g. [94]). The spectral energy density parameter $\Omega_{\rm GW}$ is related to the GW intensity, $I$, through the relation
@@
$\langle\delta n^{i}({\bf n},t)\delta n^{j}({\bf q},t^{\prime})\rangle$ $=$ $\frac{3H_{0}^{2}}{32\pi^{3}}\int df\,\frac{\bar{\Omega}_{\rm GW}(f)}{f^{3}}{\cos[2\pi f(t-t^{\prime})]}\left[H^{(0)}_{ij}({\bf n},{\bf q})+\beta\,(4-n_{\Omega})\,H^{(1)}_{ij}({\bf n},{\bf q},{\bf v})\right]\,$, (2.9)
@@
with the projector
$\mathcal{R}_{ijk}(\mathbf{n},\mathbf{p})=\frac{n_{j}}{2}\left[\frac{(n_{i}+p_{i})\,n_{k}}{1+\mathbf{n}\cdot\mathbf{p}}-\delta_{ik}\right].$ (2.12)
arising from properties of the polarization tensors in this context [25]. The tensors $H^{(0,1)}_{ij}$ are the analog of the PTA ORF relative to the monopolar and dipolar kinematic components of the SGWB. To compute these quantities, we introduce the projection tensor
@@
The integrals (2.10), (2.11), can be performed with a standard methods of contour integration used in the context of PTA physics, see e.g. [98]  [^5] In fact, the calculation proceeds exactly as in [98]. We decompose the angular integration measure as $d^{2}\Omega_{\mathbf{p}}=d(\cos\theta)\,d\varphi$, and perform the substitution $\varphi=e^{iz}$. In this way, the resulting integral in the complex plane is expressed as a contour integral over $\mathcal{C}$, corresponding to the unit circle. The integrand has poles inside the unit circle, and the integral and can be straightforwardly evaluated using Cauchy’s integral formula and the residue theorem.. For the case of the monopole, the ORF is associated with the integral in eq (2.10). We consider the combination
$y$ $=$ $\frac{1-{\bf n}\cdot{\bf q}}{2}\,=\,\frac{1-\cos{\zeta}}{2}\,$ (2.14)
@@
a result equivalent to what was found in [25], although expressed slightly more compactly. As mentioned above, we obtained this expression by evaluating the integral (2.10) using contour integration. The expression (2.15) is the astrometry equivalent of the Hellings-Downs function, the PTA ORF for the SGWB monopole. From the statistical properties of the two-point correlators, one expects a dependence only on the angular separation ${\bf n}\cdot{\bf q}$ between stars in (2.15). This feature is not immediately apparent in our results because of the tensorial structure introduced by the projection tensor (2.13), which brings explicit factors of $n_{i}$ and $q_{j}$ into the expression. A useful way to reveal this property is through the quantity ${\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]$ [^6] This expression also plays an important role in the Fisher forecasts of section 3., which, as expected, depends only on the relative directions (see Appendix B). Alternative formulations that make the exclusive dependence on ${\bf n}\cdot{\bf q}$ manifest have also been explored; see, e.g., [28, 31].
@@
${\bf A}$ $=$ ${\bf n}\times{\bf q}\qquad,\qquad{\bf B}\,=\,{\bf n}\times{\bf A}\qquad,\qquad{\bf C}\,=\,-{\bf q}\times{\bf A}\,.$ (2.16)
@@
$(Av)\,=\,{\bf v}\cdot\left({\bf n}\times{\bf q}\right)\qquad,\qquad(nv)\,=\,{\bf n}\cdot{\bf v}\,$, (2.18)
@@
Our analytical, covariant expression (2.17) for the ORF demonstrates that the astrometry sensitivity to the SGWB dipole depends not only on the angle among stars, $\zeta$, but also on their position in the sky, and on the velocity vector $\bf v$} among frames.
The compact expressions (2.15) and (2.17) are particularly suitable to visualise patterns of sensitivity on the celestial sphere, and to further explore geometrical features of physically relevant quantities in the context we are examining. We plot in Figure 1 the combinations.
@@
Figure 1: The quantities ${\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]$ and ${\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]$, associated with response to a SGWB depending on stars positions. (See the main text for our notation.) Motivated by CMB, we choose the kinematic dipole direction $\bf v$} (red star) at $(l,b)=(264^{\circ},48^{\circ})$ in galactic coordinates. Each panel shows a different choice of $\bf n$}, while the stars $\bf q$} take the position of each pixel of the map. Upper left panel: $\bf n$} pointing towards $(l,b)=(0,0)$. Upper right panel: $\bf n$} pointing towards $\bf v$}. Lower left panel: $\bf n$} at $(l,b)=(0,0)$. Lower right panel: $\bf n$} pointing towards the direction $(l,b)=(270.21^{\circ},-75.45^{\circ})$.
@@ 2.2 Cross-correlated overlap functions of astrometry and PTA
$\langle\delta n_{i}({\bf n},t)z(t)\rangle\,=\,\frac{3\,H_{0}^{2}}{64\pi^{3}}\int df\frac{\bar{\Omega}_{\rm GW}(f)}{f^{3}}\,\left[K_{i}^{(0)}({\bf n},{\bf x})+\beta\,(4-n_{\Omega})\,K_{i}^{(1)}({\bf n},{\bf x},{\bf v})\right]\,.$ (2.22)
@@
$K^{(0)}_{i}({\bf n},{\bf x})$ $=$ $\int d^{2}\Omega_{{\bf p}}\,{\cal R}_{ikl}({\bf n},{\bf p})\,P_{klrs}\,\frac{x^{r}x^{s}}{1+{\bf x}\cdot{\bf p}}\,$, (2.23)
$K^{(1)}_{i}({\bf n},{\bf x},{\bf v})$ $=$ $\int d^{2}\Omega_{{\bf p}}\,({\bf p}\cdot{\bf v})\,\,{\cal R}_{ikl}({\bf n},{\bf p})\,P_{klrs}\,\frac{x^{r}x^{s}}{1+{\bf x}\cdot{\bf p}}\,.$ (2.24)
@@
$K^{(0)}_{i}({\bf n},{\bf x})$ $=$ $\frac{16\pi}{3}\frac{(1-2y)\,n_{i}-x_{i}}{4y(1-y)}\,\left(2y-2y^{2}+3y^{2}\ln(y)\right)\,$, (2.26)
@@
We call $\zeta_{sp}$ the angle between star and pulsar directions, $\cos\zeta_{sp}={{\bf n}}\cdot{\bf{x}}$. We represent a simple example in the upper left panel of Figure 3 with $\bf{n}$ pointing towards $(l,b)=(0,0)$; the roots of the function $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$ are at $\zeta_{sp}=0,\pi$ and $\zeta_{sp}\approx 86.14^{\circ}$, while its local maxima are located at $\zeta_{sp}\approx 37.13^{\circ}$ and $\zeta_{sp}\approx 132.195^{\circ}$, and are represented in the left panel of Figure 4. The lower panel of Figure 3 shows two scenarios for the general expression for the function $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ (see also Appendix B), which depends on the angle between star and pulsar and the angle between the velocity and pulsars given by $vx={\bf v}\cdot{\bf{x}}$; on the left we choose $\bf{n}$ at $(l,b)=(0,0)$, and on the right, $\bf{n}$ is at position $(l,b)=(270.21^{\circ},-75.45^{\circ})$ (a direction chosen randomly). The sensitivity patterns are complex. However, in the specific case of the star direction $\bf{n}$ pointing towards the direction $\bf v$} we obtain a simple formula, depending only on the angle between star and pulsar. This last case is shown on the right upper panel of Figure 3, where the bluest, less sensitive regions showing the positions of objects at relative angle $\zeta_{sp}=0,\pi$ and $39.82^{\circ},101.37^{\circ}$, corresponding to minimal sensitivity to dipole. The red region shows the positions of the pulsars where the function $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ reaches its maximum at $\zeta_{sp}\approx 142.29^{\circ}$. There are also local maxima at $\zeta_{sp}\approx 17.97^{\circ},71.88^{\circ}$, which are more apparent on the right panel of Figure 4. Notice that the quantities ${\bf K}_{0,1}$ corresponding to the synergetic ORF vanish in the concident limit of the star aligned with the pulsar direction. One may understand this as follows – the star deflection always lies in the plane perpendicular to the line of sight. On the contrary, the pulsar timing residual arises from the change in length along the line of sight. Thus for pulsar and star in same direction the two effects are perpendicular to each other.
@@
Figure 3: The quantities $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$, and $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ response to stars and pulsars positions. The dipole direction $\bf v$} (red star) is chosen in the direction $(l,b)=(264^{\circ},48^{\circ})$ in galactic coordinates. Each panel shows a different choice of $\bf n$}, while the pulsars positions $\bf{x}$} scan over each pixel of the map. Upper left panel: $\bf n$} towards $(l,b)=(0,0)$. Upper right panel: $\bf n$} towards $\bf v$}. Lower left panel: $\bf n$} at $(l,b)=(0,0)$. Lower right panel: $\bf n$} towards $(l,b)=(270.21^{\circ},-75.45^{\circ})$.
@@
Figure 4: Overlap functions response in terms of the angle between the stars and pulsars. Left: Angular dependence for $\mathbf{K}_{0}\mathbf{K}_{0}^{T}$. Right: The quantities $\mathbf{K}_{1}\mathbf{K}_{1}^{T}$ when choosing ${\bf n}=-{\bf v}$.
3 Fisher forecasts
@@
In this section we use the Fisher matrix approach to forecast the sensitivity of astrometry, individually and jointly with PTA, to the SGWB properties such as the amplitude, spectral index and dipole anisotropy. In appropriate limits and special cases, the Fisher matrix obeys interesting scaling relations which make our analysis particularly simple. In general though, numerical work is needed to evaluate at what extent the synergy between astrometry and PTA improve the sensitivity to SGWB properties. We carry out such analysis in sections 3.1, 3.2 and 3.2.3, making use of the weak signal approximation for astrometry to simplify the numerical calculations for a large number of stars.
3.1 Astrometry only
We start by discussing the case of astrometry only. We expand the deflection correlation of eq (2.9) in a basis with coefficients $p_{n}$, [^9] In principle, these basis coefficients can also be spherical harmonics coefficients or individual pixels used to discretize the SGWB intensity map. Our analysis represents a special case where we only expand the SGWB intensity up to the dipole term, assuming a known dipole direction.
$\langle{\delta n_{a}^{i}\,\delta n_{b}^{j}}\rangle\equiv C^{ij}_{ab}=\sum_{n=0,1}p_{n}H_{ab,n}^{ij}+N_{ab}^{ij}$ (3.1)
The indexes $a,b$ identify the two stars. $i,j$ the three-dimensional vector components, and the index $n$ runs over $0,1$ indicating monopolar and dipolar contributions to the signal. We include the effects of noise through the noise matrix $N_{ab}^{ij}$. The quantities $p_{n}$ and $H_{ab,n}$ are controlled by the properties of the SGWB, as well as the astrometry ORF. See equation (2.9) and the discussion that follows for the definitions. The angle brackets on the LHS denote the expectation value of the correlation of the $i$-th deflection of star $a$ with the $j$-th deflection of star $b$. To keep our notation compact we now use bold symbols to represent the signal covariance as well as the noise matrix, i.e.
$C^{ij}_{ab}\equiv{\boldsymbol{C}}=\sum_{n=0,1}p_{n}H_{ab,n}^{ij}+N_{ab}^{ij}\equiv\sum_{n=0,1}p_{n}{\boldsymbol{H}}_{n}+{\boldsymbol{N}}\,.$ (3.2)
Thus, in the following the trace of the product of two bold quantities represents a sum over the $i,j$ and $a,b$ indices, i.e.
${\rm Tr}[\boldsymbol{A}\boldsymbol{B}]=\sum_{ij,ab}A^{ij}_{ab}B_{ab}^{ij}\,.$ (3.3)
At this stage, we do not need to specify whether we work either in time or in frequency domain.
Our aim is to determine the how well we can measure the quantities $p_{0,1}$, and extract from this information the properties of the SGWB: its monopole amplitude $\Omega_{\rm GW}$ as well as the value of the parameter $\beta$ controlling kinematic anisotropies, using a Fisher matrix approach. [^10] We also discuss the optimal estimators for the SGWB monopole and dipole in appendix C. However, before proceeding to build the Fisher matrix for the monopole and dipole amplitudes, we note that the matrix $C$ can not be directly be used as a covariance matrix, since it is singular. This stems from the fact that the three Cartesian deflection components are not independent since observations get projected on to the two-dimensional celestial sphere. Hence, the star deflections can be entirely described in terms of two angles $\delta\theta$ and $\delta\phi$ [28]. Let $P$ be the matrix that converts the deflections from 3D Cartesian to 2D polar coordinates, i.e.